OpenStudy (anonymous):

Show that x+y is a factor of x^(2n-1)+y^(2n-1) for all natural numbers n. Use mathematical induction.

OpenStudy (anonymous):

Why do we have to do these by induction? =/ we could just use the binomial theorem and be done in a line or two.

OpenStudy (anonymous):

My professor requires all of them to be solved with induction :\

OpenStudy (anonymous):

I'll assume you can do the Base Step in the inductive process, so i'm going to start with the Inductive step (assuming the statement is true for n = k, and using this to prove the statement is true for n = k+1)

OpenStudy (anonymous):

is that ok? or do we need to go over the Base Step too?

OpenStudy (anonymous):

That's fine. I have the base step done.

OpenStudy (anonymous):

if n = 0 does it hold?

OpenStudy (anonymous):

ok, So lets assume for some fixed natural number k that: \[(x+y)\mid x^{2n-1}-y^{2n-1}\] We want to show that: \[(x+y)\mid x^{2n+1}-y^{2n+1}\]

OpenStudy (anonymous):

That notation\[a\mid b\]means "a divides b" btw, i dont know if thats clear or not.

OpenStudy (anonymous):

when n = 0 \[\frac{1}{x}+\frac{1}{y} = \frac{x + y}{xy}\]

OpenStudy (anonymous):

Using our Inductive Hypothesis, we have:\[(x+y)\mid x^{2k-1}-y^{2k-1}\iff x^{2k-1}-y^{2k-1}=b(x+y)\]for some integer b. Solving for x^(2k-1) gives: \[x^{2k-1}=b(x+y)+y^{2k-1}\] Now looking at the statement:\[x^{2k+1}-y^{2k+1}\]we can rewrite it as:\[x^2\cdot x^{2k-1}-y^{2k+1}\]

OpenStudy (anonymous):

Now, we are going to take our Inductive Hypothesis and substitute it into that statement.

OpenStudy (anonymous):

So that nets us: \[x^2(b(x+y)+y^{2k-1})+y^{2k+1}\] Now, the name of the game is to see if we can manipulate this statement to become something like: \[(x+y)(junk)\] the junk part is going to look ugly, but if we can pull out a (x+y) like that we are good.

OpenStudy (anonymous):

\[\frac{b(x+y)*x^2y+x^2y^{2k}+y^{2k+2}}{y}\]

OpenStudy (anonymous):

i dont see it yet >.< do you Satellite?

OpenStudy (anonymous):

How can we get that x+y out of there

OpenStudy (anonymous):

oh i wasn't looking. just trying to recall how to prove all horses are the same color by induction. i know how to prove you will not be hanged by induction

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

jailer says you will be hung before saturday but you will not know the day in advance. prisoner says ok well if i will not know the day in advance then they cannot wait until saturday because that is the last day, and i will know on friday that they will hang me saturday. but if they cannot hang me on saturday surely they cannot hang me on friday, because since i will know the day before, and i can't be hung on saturday, then i cannot be hung on friday. etc.

OpenStudy (anonymous):

oh im so stupid >.< its because its late. I made a sign error typing it out. it should be: \[x^{2k+1}-y^{2k+1}=x^2(x^{2k-1})-y^{2k+1}\]\[=x^2(b(x+y)+y^{2k-1})-y^{2k+1}=bx^2(x+y)+x^2y^{2k-1}-y^{2k+1}\]\[bx^2(x+y)+y^{2k-1}(x^2-y^2)=bx^2(x+y)+y^{2k-1}(x+y)(x-y)\]

OpenStudy (anonymous):

i think the gimmick is to factor out and subtract off. something like \[x^{2n+1}+y^{2n+1}=(x^{2n-1}+y^{2n-1})x^2y^2-stuff\]

OpenStudy (anonymous):

Now you can factor out the (x+y) and you get: \[(x+y)(bx^2+y^{2k-1}(x-y))\] This concludes my Proof by Mathematical Induction. (stupid sign errors >.>)

OpenStudy (akshay_budhkar):

Well done professor Joesph!!!!! and hi joe

OpenStudy (anonymous):

yo :)

OpenStudy (anonymous):

Satellite! i need you help with a homework problem lol ima post a new question.

OpenStudy (anonymous):

im gonna scan it, one sec.

OpenStudy (anonymous):

ok but it may have to wait until the morning. let me see i can do this one

OpenStudy (akshay_budhkar):

professor needs help from LEGAND COOL

OpenStudy (anonymous):

\[x^{2n+1}-y^{2n+1}=(x^{2n-1}-y^{2n-1})x^2y^2-stuff\]

OpenStudy (anonymous):

If all else fails... remember to use logarithms.

OpenStudy (anonymous):

now what is stuff?

OpenStudy (anonymous):

im gonna post a new question., posting....now <.<

OpenStudy (anonymous):

ah i can't do this it is too late. i will look at yours in the morning

OpenStudy (anonymous):

lol fine <.<

OpenStudy (nikvist):

\[x^{2k+1}+y^{2k+1}=\]\[=x^{2k+1}+x^2y^{2k-1}-x^2y^{2k-1}-x^{2k-1}y^2+x^{2k-1}y^2+y^{2k+1}=\] \[=x^2(x^{2k-1}+y^{2k-1})-x^2y^2(x^{2k-3}+y^{2k-3})+y^2(x^{2k-1}+y^{2k-1})=\] \[=(x^2+y^2)(x^{2k-1}+y^{2k-1})-x^2y^2(x^{2k-3}+y^{2k-3})\equiv 0\,\,\,(mod\,\,(x+y))\]